# X-Portrait: Expressive Portrait Animation with Hierarchical Motion Attention

### v3

@@ 3.3 Enhanced Attention to Local Facial Movements
Our trained control module $\mathcal{C}$ offer a significant improvement over coarse landmarks in capturing head movements and low-frequency facial expressions. However, during inference, we occasionally observe our control module fails to capture nuanced motion exhibited in the driving video, such as subtle nose wrinkles and intricate eyeball movements. These fine-grained motions, often overlooked by prior works, are crucial for conveying emotions and overcoming the uncanny valley artifacts in portrait animation.
@@
Our control module $\mathcal{C}$, formulated as a ControlNet [50], extracts structural features from the conditional image $I_{C}$ and integrates into the UNets via skip connections during the denoising process. However, such additive conditional attention operates in the global image space, treating motion in every pixel with equal weight. This observation has motivated us to guide the control module to enhance localized attention specifically to critical facial regions, aimed at better animation realism and finer control granularity.
@@
We achieve motion control at nuanced scales by introducing an auxiliary ControlNet that conditions on a masked image $I^{l}_{C}$ revealing only patches around the eyes and mouth from $I_{C}$. Specifically, we detect landmarks of the control image for the eyes and mouth, using their centers to crop $128\times 128$ patches as local control signals $I^{l}_{C}$. While simple, this control branch effectively provides enhanced guidance to the UNet denoising, focusing solely on the local structure extracted from those cropped facial regions. Notably, before the training, we follow the same philosophy to fine tune the control image generator $\mathcal{F}$ with additional L1 and VGG [22] perceptual reconstruction losses on the local patches $I^{l}_{C}$. Our enhanced generator, named as Face Vid2Vid Plus, helps capturing the subtle motions in the hierarchical conditional inputs $I_{C}$ and $I^{l}_{C},$ benefiting the subsequent training of both control modules.
3.4 Identity Preservation
@@
Figure 4: (a) is the result of X-Portrait  with a single reference image from (top, b), while (c) uses both references. X-Portrait  seamlessly accommodates multiple reference images, producing animations with better captured personalized appearance traits in (c). Please find the differences in hair, ear and face shape.
@@
Figure 5: Qualitative comparisons. Among all the methods, X-Portrait  achieves the most accurate and robust transfer of both subtle and extreme facial expressions (e.g., pouting and single-eye blinks) and wide-range head translations and rotations, with precise identity resemblance (e.g., face shapes, eyes/mouth sizes) to the reference even with artistic styles (e.g., pencil drawing and anime).
With a single reference source image $I_{S},$ only partial facial appearance is visible, and the network has to rely on the universal generative prior of LDM for inpainting unobserved facial regions when altering head poses or camera views. However, when more reference images are accessible, such as in a video, we can incorporate a more comprehensive appearance context without any network modification. Owing to our disentangled controls, by simply concatenating the multiple extracted appearance features into the UNets with $\mathcal{R}$, our framework can seamlessly fuse them and generate animations with better-retained identity attributes (Figure 4).
4 Experiments
4.1 Implementation Details
@@
We train our model using an in-house dataset including monocular camera recordings of 42 expressions and 20-min talkings from 550 subjects in both indoor and outdoor scenes, with a cropped resolution of $512\times 512$. For data processing, we follow the approach outlined in FOMM [35] but enlarge the cropping region to include the shoulders. For evaluation, we collect 100 in-the-wild portraits [29, 9, 28] in various realistic or artistic depictions (2D/3D cartoon, anime, cyberpunk, oil painting, statue, wood, etc.), facial appearances (joker, elf, human-like robot, etc.), apparels (glasses, hat, robe, headphones etc.), and body poses (front and side). Additionally, we collect 200 licensed test videos with varying emotions, head poses, and facial expressions.
@@
We utilize Stable Diffusion 1.5 as our generative backbone and we keep its weights frozen during the entire training phases. Our training was conducted in stages, where we sequentially plug in and train the modules $\mathcal{R},\mathcal{C}$ and $\mathcal{M}$ With 8 A100s, each module undergoes training with $30K$ steps with 16 video frames in each step. During inference, we leverage the prompt traveling strategy [42] to enhance temporal smoothness. With a frozen UNet, X-Portrait demonstrates inherent compatibility with the latent consistency model [27]. This compatibility facilitates the efficient generation of a 24-frame animation within 30 seconds (10 steps) when executed on an A10 GPU. Notably, instead of denoising from random Gaussian noise, we apply the forward diffusion process on source image $I_{S}$ into an initialized noise. Such generated noise adds a subtle level of structural guidance at the early denoising step, yielding improved consistency with reduced popping artifacts. We do not rely on $\mathcal{F}$ during inference.
Table 1: Quantitative comparisons of X-Portrait with SOTA baselines in self and cross reenactments, evaluated on $256\times 256$ test images.
@@
Face Vid2vid [43] 0.042 0.770 0.203 52.071 0.615 36.222 0.100/4.72
DaGAN [20] 0.054 0.732 0.234 64.297 0.280 35.537 0.111/5.10
TPS [51] 0.038 0.801 0.179 50.636 0.508 34.904 0.079/3.94
MCNet [19] 0.035 0.811 0.176 51.091 0.343 34.624 0.094/3.82
FADM [47] 0.049 0.727 0.231 51.091 0.635 34.640 0.103/4.62
MagicDance [6] 0.045 0.766 0.172 30.383 0.643 67.241 0.099/7.04
@@
4.2 Evaluations and Comparisons
Our method empowers the creation of captivating and highly expressive animations, demonstrating a diverse range of head motions (with rotations over 150 degrees) and facial expressions (frowning, crossed eyes, pouting etc) across both realistic, human-like, and style portraits (Figure 178). Note that X-Portrait employs a reference module to effectively cross-query source appearance features, thereby establishing localized spatial correspondences between the input and output. Once trained, the model is able to generalize to out-of-domain appearances through its learned latent space, as exemplified by stylized portraits. Simultaneously, we maintain high identity resemblance to the given source image throughout the generated video. We compare our method with prior portrait animation works including state-of-the-art GAN based methods Face Vid2vid [43] (and our enhanced version Face Vid2vid Plus), TPS [51], DaGAN [20], MCNet [19] and recent diffusion-based approaches like FADM [47] and MagicDance [6]. We exclude MegaPortraits [10] from our comparisons as there is no public release. For fair comparisons, we fine-tune all the baselines over the same dataset. Notably, for MagicDance [6], we use our own implementation and fine-tune the 68-landmark-based ControlNet from the public checkpoint. We assess their performances over both self and cross reenactments. All numbers are computed at the resolution of $256\times 256$ due to the limited resolution for most of the previous works.
@@ 4.2 Evaluations and Comparisons
As evidenced in our qualitative comparisons in Figure. 5, X-Portrait  demonstrates the best perceptual quality, motion precision and identity similarity. Despite being trained exclusively on real portrait animation, our method exhibits superior domain generalization capabilities to style portraits, whereas all GAN-based baselines suffer from severe blurriness for out-of-domain styles and under large head motions. For both extreme (e.g., pouting, top rows) and subtle expressions (e.g., single eye blinks, bottom rows), all the methods except ours fail to capture and transfer the nuances to the reference portraits. MagicDance [6] yields unstable animations when landmarks detection fails (middle rows) and identity drifts when the source and driving have distinct facial features (bottom rows).
For quantitative assessment, given the absence of image ground truth, we employ three metrics to evaluate identity similarity, image quality, and expression and head pose accuracy, respectively. Specifically, the ArcFace score [7], calculating the cosine similarity between source and generated images, is utilized to assess identity preservation. We employ a pre-trained network (HyperIQA [39]) for image quality assessment. To evaluate the motion accuracy, we compare the L1 difference between the extracted facial blendshapes and head poses of the driving and generated frames using ARKit [2]. As reported in Table 1, our method consistently outperforms all competitors by a good margin. Furthermore, a user study was carried out, revealing that an overwhelming majority of participants, comprising approximately $83.23\%$, affirmed that X-Portrait exceeds FaceVid2Vid Plus in terms of motion accuracy and expressiveness. Notably, by leveraging the Stable Diffusion prior, X-Portrait  and MagicDance surpasses the other methods by a substantial margin in image quality. However, compared to MagicDance, X-Portrait  shows superior identity preservation and motion accuracy, thanks to our novel hierarchical cross-identity scale-augmented controls.
Table 2: Quantitative ablation.
@@ 4.3 Ablation Studies
We ablate the efficacy of individual components by removing them from our full training pipeline, evaluated on cross reenactment synthesis. Figure 3 (a) demonstrates the results when training the model naively with the driving frame as both the target and motion condition (self-driven training, even with our scaling strategy). In this scenario, the network tends to treat it as an image reconstruction task and merely copies both the identity and motion from the driving frames. Therefore, as shown in the quantitative evaluation(Table 2 (a)), while the expression accuracy is on par with our full pipeline, there is a significant decrease in identity resemblance. Figure 3 (b) highlights the importance of our local control module in capturing subtle driving expressions. Excluding the local control module results in the absence of expression details, such as the asymmetric frowning, aligning with the observation of decreased expression accuracy (Table 2 (b)). We note that it achieves a slightly better identity similarity score due to the more rigid local facial movements. Furthermore, the source identity features are better maintained with our scaling augmented training strategy without which noticeable identity drift to the driving occurs, as evidenced in Figure 3 (c) and Table 2 (c). Lastly, we qualitatively demonstrate in Figure 4 that personal identity traits can be enhanced and seamlessly fused into our framework with multiple reference images.
@@
Figure 6: X-Portrait exhibits limited transfer of expression when $\mathcal{F}$ completely fails to produce any correlated motion clues (e.g., lips inward (top) and puffing cheeks (bottom)).
4.4 Limitations and Future Work
In the future, we aim to enhance the model’s capabilities by animating gestures to further improve expressiveness. As a parallel objective, we recognize that the synthesized image quality could benefit from refined base diffusion models, especially in the teeth region. Additionally we would like to explore more advanced spatiotemporal attentions for smooth video generation without noticeable jittering artifacts. X-Portrait  represents a general pipeline that improves the expressiveness of the base reenactment method, i.e., $\mathcal{F}$. However, as illustrated in Figure. 6, our method might be limited in depicting some extremely challenging expressions especially when $\mathcal{F}$ completely fails to produce any motion clues.
5 Conclusion
@@
We introduce X-Portrait, a novel stable diffusion-based framework crafted for portrait animation, ensuring meticulous transfer of driving facial expressions and head poses. Our method excels with the incorporation of cross-identity driving inputs in training, facilitating a balanced achievement of motion expressiveness, identity preservation, and animation robustness. Additionally, we propose a local control module to accentuate the attention to detailed facial expressions that are subtle to capture but critical to emotion conveyance. The showcased impressive performance of our model on generalized source portraits and driving motions validates its effectiveness.
@@
Figure 7: More qualitative results.The leftmost column shows the driving frame, and we show the animated results for multiple portraits by X-Portrait.
@@
Figure 8: More qualitative results.The leftmost column shows the driving frame, and we show the animated results for multiple portraits by X-Portrait.

### v4

@@ 3.3. Enhanced Attention to Local Facial Movements
Our trained control module $\mathcal{C}$ offer a significant improvement over coarse landmarks in capturing head transformations and low-frequency facial expressions. However, during inference, we occasionally observe our control module fails to capture nuanced motion exhibited in the driving video, such as subtle nose wrinkles and intricate eyeball movements. These fine-grained motions, often overlooked by prior works, are crucial for conveying emotions and overcoming the uncanny valley artifacts in portrait animation.
@@
Our control module $\mathcal{C}$, formulated as a ControlNet (Zhang et al., 2023a), extracts structural features from the conditional image $I_{C}$ and integrates into the UNets via skip connections during the denoising process. However, such additive conditional attention operates in the global image space, treating motion in every pixel with equal weight. This observation has motivated us to guide the control module to enhance localized attention specifically to critical facial regions, aimed at better animation realism and finer control granularity.
@@
We achieve motion control at nuanced scales by introducing an auxiliary ControlNet that conditions on a masked image $I^{l}_{C}$ revealing only patches around the eyes and mouth from $I_{C}$. Specifically, we detect landmarks of the control image for the eyes and mouth, using their centers to crop patches of $128\times 128$ as local control signals $I^{l}_{C}$. While simple, this control branch effectively provides enhanced guidance to the UNet denoising, focusing solely on the local structure extracted from those cropped facial regions. Notably, before the training, we follow the same philosophy to fine tune the control image generator $\mathcal{F}$ with additional L1 and VGG (Johnson et al., 2016) perceptual reconstruction losses on the local patches $I^{l}_{C}$. Our enhanced generator, named as Face Vid2Vid Plus, helps capturing the subtle motions in the hierarchical conditional inputs $I_{C}$ and $I^{l}_{C},$ benefiting the subsequent training of both control modules.
3.4. Identity Preservation
@@
Figure 4. Qualitative comparisons. Among all the methods, X-Portrait  achieves the most accurate and robust transfer of both subtle and extreme facial expressions (e.g., pouting and single-eye blinks) and wide-range head translations and rotations, with precise identity resemblance (e.g., face shapes, eyes/mouth sizes) to the reference even with artistic styles (e.g., pencil drawing and anime). ⓒArtbyhoussam, Midjourney.com and Katiin Bolovtsova.
With a single reference source image $I_{S},$ only partial facial appearance is visible, and the network has to rely on the universal generative prior of LDM for inpainting unobserved facial regions when altering head poses or camera views. However, when more reference images are accessible, such as in a video, we can incorporate a more comprehensive appearance context without any network modification. Owing to our disentangled controls, by simply concatenating the multiple extracted appearance features into the UNets with $\mathcal{R}$, our framework can seamlessly fuse them and generate animations with better-retained identity attributes (more details in Supplement Section A).
4. Experiments
4.1. Implementation Details
@@
We train our model using an in-house dataset including monocular camera recordings of 42 expressions and 20-min talkings from 550 subjects in both indoor and outdoor scenes. All the data were processed with a cropped resolution of $512\times 512$. Sequences of low quality were filtered out with (Hosu et al., 2020). All videos feature real subjects showcasing a diverse range of expressions and speeches in various scenes. For data processing, we follow the approach outlined in FOMM (Siarohin et al., 2019b) but enlarge the cropping region to include the shoulders. For evaluation, we collect 100 in-the-wild portraits (Pexels, 2024; DeviantArt, 2024; Midjourney, 2024) in various realistic or artistic depictions (2D/3D cartoon, anime, cyberpunk, oil painting, statue, wood, etc.), facial appearances (joker, elf, human-like robot, etc.), apparels (glasses, hat, robe, headphones etc.), and body poses (front and side). Additionally, we collect 200 licensed test videos with varying emotions, head poses, and facial expressions.
@@
We utilize SD 1.5 as our generative backbone and we freeze its weights during the entire training phases. Prior to training, the control and appearance reference modules are initialized using SD 1.5, whereas the motion module is initialized with the weights of AnimateDiff (Guo et al., 2023b). Our training was conducted in stages, where we sequentially plug in and train the modules $\mathcal{R},\mathcal{C}$ and $\mathcal{M}$ An AdamW optimizer is utilized with a learning rate of $10^{-5}$ to train all modules. Each module undergoes training with $30K$ steps with 16 video frames in each step.
During inference, we leverage the prompt traveling strategy (Tseng et al., 2022) to enhance temporal smoothness. With a frozen SD UNet, X-Portrait demonstrates inherent compatibility with the latent consistency model (Luo et al., 2023). This compatibility facilitates the efficient generation of a 24-frame animation within 30 seconds (10 steps) when executed on an A10 GPU. Notably, instead of denoising from random Gaussian noise, we apply the forward diffusion process on source image $I_{S}$ into an initialized noise. Such generated noise adds a subtle level of structural guidance at the early denoising step, yielding improved consistency with reduced popping artifacts. We do not rely on $\mathcal{F}$ during inference.
Table 1. Quantitative comparisons of X-Portrait  with SOTA baselines in self and cross reenactment tasks, evaluated on an image resolution of $256\times 256$
@@
Face Vid2vid (Wang et al., 2021) 0.042 0.770 0.203 52.071 0.615 36.222 0.100/4.72
DaGAN (Hong et al., 2022) 0.054 0.732 0.234 64.297 0.280 35.537 0.111/5.10
TPS (Zhao and Zhang, 2022) 0.038 0.801 0.179 50.636 0.508 34.904 0.079/3.94
MCNet (Hong and Xu, 2023) 0.035 0.811 0.176 51.091 0.343 34.624 0.094/3.82
FADM (Zeng et al., 2023) 0.049 0.727 0.231 51.091 0.635 34.640 0.103/4.62
MagicDance (Chang et al., 2024) 0.045 0.766 0.172 30.383 0.643 67.241 0.099/7.04
@@
4.2. Evaluations and Comparisons
Our method empowers the creation of captivating and highly expressive animations, demonstrating a diverse range of head motions (with rotations over 150 degrees) and facial expressions (frowning, crossed eyes, pouting etc) across both realistic, human-like, and style portraits (Figure 167). Note that X-Portrait employs a reference module to effectively cross-query source appearance features, thereby establishing localized spatial correspondences between the input and output. Once trained, the model is able to generalize to out-of-domain appearances through its learned latent space, as exemplified by stylized portraits. Simultaneously, we maintain high identity resemblance to the given source image throughout the generated video. We compare our method with prior portrait animation works including state-of-the-art GAN based methods Face Vid2vid (Wang et al., 2021) (and our enhanced version Face Vid2vid Plus), TPS (Zhao and Zhang, 2022), DaGAN (Hong et al., 2022), MCNet (Hong and Xu, 2023) and recent diffusion-based approaches like FADM (Zeng et al., 2023) and MagicDance (Chang et al., 2024). We exclude MegaPortraits (Drobyshev et al., 2022) from our comparisons as there is no public release. For fair comparisons, we fine-tune all the baselines over the same dataset. Notably, for MagicDance (Chang et al., 2024), we use our own implementation and fine-tune the 68-landmark-based ControlNet from the public checkpoint. We assess their performances over both self and cross reenactments. All numbers are computed at the resolution of $256\times 256$ due to the limited resolution for most of the previous works.
@@ 4.2. Evaluations and Comparisons
As evidenced in our qualitative comparisons in Figure 4, X-Portrait  demonstrates the best perceptual quality, motion precision and identity similarity. Despite being trained exclusively on real portrait animation, our method exhibits superior domain generalization capabilities to style portraits, whereas all GAN-based baselines suffer from severe blurriness for out-of-domain styles and under large head motions. For both extreme (e.g., pouting, top rows) and subtle expressions (e.g., single eye blinks, bottom rows), all the methods except ours fail to capture and transfer the nuances to the reference portraits. MagicDance (Chang et al., 2024) yields unstable animations when landmarks detection fails (middle rows) and identity drifts when the source and driving have distinct facial features (bottom rows).
For quantitative assessment, given the absence of image ground truth, we employ three metrics to evaluate identity similarity, image quality, and expression and head pose accuracy, respectively. Specifically, the ArcFace score (Deng et al., 2019), calculating the cosine similarity between source and generated images, is utilized to assess identity preservation. We employ a pre-trained network (HyperIQA (Su et al., 2020)) for image quality assessment. To evaluate the motion accuracy, we compare the L1 difference between the extracted facial blendshapes and head poses of the driving and generated frames using ARKit (Apple, 2023). As reported in Table 1, our method consistently outperforms all competitors by a good margin. Furthermore, a user study was carried out to affirm that X-Portrait significantly exceeds FaceVid2Vid Plus in terms of motion accuracy and expressiveness (more details in Supplement Section C). Notably, by leveraging the SD prior, X-Portrait  and MagicDance surpasses the other methods by a substantial margin in image quality. However, compared to MagicDance, X-Portrait  shows superior identity preservation and motion accuracy, thanks to our novel hierarchical cross-identity scale-augmented controls.
Table 2. Quantitative ablation.
@@ 4.3. Ablation Studies
We ablate the efficacy of individual components by removing them from our full training pipeline, evaluated on cross reenactment synthesis. Figure 3 (a) demonstrates the results when training the model naively with the driving frame as both the target and motion condition (self-driven training, even with our scaling strategy). In this scenario, the network tends to treat it as an image reconstruction task and merely copies both the identity and motion from the driving frames. Therefore, as shown in the quantitative evaluation(Table 2 (a)), while the expression accuracy is on par with our full pipeline, there is a significant decrease in identity resemblance. Figure 3 (b) highlights the importance of our local control module in capturing subtle driving expressions. Excluding the local control module results in the absence of expression details, such as the asymmetric frowning, aligning with the observation of decreased expression accuracy (Table 2 (b)). We note that it achieves a slightly better identity similarity score due to the more rigid local facial movements. Furthermore, the source identity features are better maintained with our scaling augmented training strategy without which noticeable identity drift to the driving occurs, as evidenced in Figure 3 (c) and Table 2 (c).
4.4. Limitations and Future Work
X-Portrait  represents a general pipeline that improves the expressiveness of the base reenactment method $\mathcal{F}$. Nevertheless, due to representation limitation, $\mathcal{F}$ may not activate at all for cases with exceptionally challenging expressions. In such instances, these expressions are filtered out by $\mathcal{F}$ during training, posing challenges for X-Portrait to learn the corresponding distribution, as illustrated in Figure 5.
@@ 4.4. Limitations and Future Work
Figure 5. X-Portrait exhibits limited transfer of expressions when $\mathcal{F}$ completely fails to produce any correlated motion clues (e.g., turning lips inwards in the first row and puffing cheeks in the second row).
In the future, we aim to enhance the model’s capabilities by animating gestures to further improve expressiveness. As a parallel objective, results quality could benefit from refined base diffusion models, e.g., the teeth region, and advanced spatiotemporal attentions could be explored to reduce noticeable jittering artifacts.
5. Conclusion
@@
We introduce X-Portrait, a novel SD-based framework crafted for portrait animation, ensuring meticulous transfer of driving facial expressions and head poses. Our method excels with the incorporation of cross-identity driving inputs in training, facilitating a balanced achievement of motion expressiveness, identity preservation, and animation robustness. Additionally, we propose a local control module to accentuate the attention to detailed facial expressions that are subtle to capture but critical to emotion conveyance. The showcased impressive performance of our model on generalized source portraits and driving motions validates its effectiveness.
@@
Figure 6. More qualitative results.The leftmost column shows the driving frame, and we show the animated results for multiple portraits by X-Portrait. ⓒMart Production, Ketut Subiyanto, Alena Darmel, Gifing.com and Cottonbro Studio.
@@
Figure 7. More qualitative results.The leftmost column shows the driving frame, and we show the animated results for multiple portraits by X-Portrait. ⓒKetut Subiyanto, Tima Miroshnichenko, Karolina Grabowska, Alex Green, Katrin Bolovtsova.